Amythehedgehog:

How do you solve ( x − 3 ) ( x + 4 ) 5 − 14 x + 35 6 = ( 52 x + 5 ) 10 ?

Ultrilliam:

oh boy... lemme make this more human readable...

pandasurvive:

It got weird :/

Amythehedgehog:

oops let me try again

Ultrilliam:

(x−3)(x+4)5−14x+356=(52x+5)10

Ultrilliam:

is that what you meant?

Amythehedgehog:

yes

Ultrilliam:

just take a close look at it and make sure it's correct, because i feel like something may be different .-.

Ultrilliam:

oh and, Welcome to QC! :)

Amythehedgehog:

Oh thank you! and let me check my problem

Amythehedgehog:

um I think some of them are fractions

Ultrilliam:

could you screenshot it?

Amythehedgehog:

how do I screen shot things

pandasurvive:

http://app.prntscr.com/en/download.html

Eiwoh2:

Hello Amythehedgehog. Welcome to QuestionCove

Ultrilliam:

Are you using windows?

Amythehedgehog:

um I am not allowed to download things, and I am very pleased to meet all of you

Amythehedgehog:

Like google Chrome?

Ultrilliam:

if so you can do this. http://prntscr.com/f5ibw6 http://prntscr.com/f5ibzr then select the area, save the image, and then upload that image here with the attach file button

KjSaif:

Simplify both sides of the equation. 5x^2 − 9x + 296 = 520x + 50


Subtract 520 x + 50 from both sides. 5x^2 − 9x + 296 − (520x + 50) = 520x + 50 − (520x + 50) ^^^^----> 5x^2 − 529x + 246 = 0
Use quadratic formula with a=5, b=-529, c=246. x= −b ± √b2−4ac / 2a x= −(−529) ± √(−529) 2 − 4(5)(246) / 2(5) x= 529 ± √274921 / 10 x= 529 / 10 + 1 / 10 √274921 or x = 529/10 + −1/10 √274921 ouch....my hands.....;-;

Amythehedgehog:

where do I use the snipping tool, I am not allowed to download things and I dont have a user to snapshot.

callielovesyhuu:

.This might be wrong but I think that the answer is y=5x²+511x+246

Ultrilliam:

Are you using windows amy?

Amythehedgehog:

Like Google Chrome Yes

Ultrilliam:

Wait, are you using Windows, or Chrome OS? O_o Your sending me mixed signals here...

Amythehedgehog:

Chrome

KjSaif:

\[\frac{ 529}{ 10} + \frac{ 1 }{ 10 } ­­ √274921\]

KjSaif:

^^answer

Amythehedgehog:

my computer is Windows

Amythehedgehog:

but I am using chrome

Ultrilliam:

Ok, do you have the clipping tool open?

Amythehedgehog:

clipping tool?

Ultrilliam:

snipping tool*

KjSaif:

*snipping

Amythehedgehog:

where is the snipping tool

KjSaif:

press on start at your bottom left

Amythehedgehog:

ok

KjSaif:

click on search programs and files

KjSaif:

and type " snipping tool"

Ultrilliam:

Wait, are you using windows 7 or windows 10?

Amythehedgehog:

one sec I got it

Amythehedgehog:

how do I paste it

Amythehedgehog:

hello?

pandasurvive:

You must save it first.

Amythehedgehog:

oh ok

pandasurvive:

Then click attach file then select the file.

Amythehedgehog:

1 attachment
Amythehedgehog:

how do you guys get profile changed?

Amythehedgehog:

how do I tag people? and how do I get my profile changed?

Amythehedgehog:

does anyone have the answer to my question?

pandasurvive:

Change your profile pic?

Amythehedgehog:

Yeah how do I do it?

pandasurvive:

Click the dolphin pic in the top right corner then you would have to select a file

Amythehedgehog:

ok

pandasurvive:

As for the tagging people just put @ then type the name like @Amythehedgehog

Amythehedgehog:

ok

Amythehedgehog:

um it wont let me change my profile

Amythehedgehog:

why is the profile thing not working?

Amythehedgehog:

O-O

Elsa213:

Gimme your password and I'll change it for you o;

Amythehedgehog:

O-O isnt that giving someone personal info?

Ultrilliam:

bruh elsa...

Elsa213:

Step 1: http://prntscr.com/f5is1l Step 2: Choose a picture Step 3: Congratulate yourself

Ultrilliam:

elsa beat me to it lol

Amythehedgehog:

um can someone tell me how to change my profile

pandasurvive:

I told you above......

Amythehedgehog:

it wont work

Elsa213:

https://questioncove.com/settings

pandasurvive:

You have to have the picture you want downloaded

Amythehedgehog:

I know I am trying but it says verify email and I keep doing it but its not working

pandasurvive:

Weird, did you verify?

Amythehedgehog:

yeah but it keeps saying verify email

Elsa213:

@Ultrilliam o:

Amythehedgehog:

@Ultrilliam

Elsa213:

I tagged him first ;o

Ultrilliam:

refresh

Ultrilliam:

WAIT

Ultrilliam:

you haven't verified your email ._.

Ultrilliam:

you may want to hit "resend" at the top and try again :P

Amythehedgehog:

how? I keep Verifying it but it keeps saying I haven't done it what is happening?

Ultrilliam:

Wait, I noticed your using a school email... Are you actually receiving the email for the verification?

Amythehedgehog:

I only recived one

Ultrilliam:

did you click the verification link?

Amythehedgehog:

verifacation link? one sec

Amythehedgehog:

um Now it says my connection is not private and I cant go back on Gmail, its blocked just great!

Ultrilliam:

O_o

Ultrilliam:

YOLO, boom your verified.

Amythehedgehog:

am I? YAY!